OpenStudy (anonymous):

sin(105 degrees) - sin(15 degrees) and sin(52.5 degrees)cos(7.5 degrees)

OpenStudy (anonymous):

\[\sin C-\sin D=2\cos \frac{ C+D }{ 2 }\sin \frac{ C-D }{2}\] \[2\sin C \cos D=\sin \left( C+D \right)+\sin \left( C-D \right)\] Now you can solve the two questions.